OpenStudy (anonymous):

(cotx - 1)(squareroot of 3 cotx + 1) = 0

OpenStudy (anonymous):

so either \(\cot(x)=1\iff \sin(x)=\cos(x)\) or \[\cot(x)=-\frac{1}{\sqrt{3}}\iff \tan(x)=\sqrt{3}\] and your job is to solve for \(x\) now it should be somewhat easier

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

whoa except i made a mistake last line should read \[\tan(x)=-\sqrt{3}\] i dropped the minus sign by mistake

OpenStudy (anonymous):

tnx again

OpenStudy (anonymous):

which makes \(x=\frac{2\pi}{3},\frac{5\pi}{3},...\) i hope you have an interval over which to solve

OpenStudy (anonymous):

what does the arrow mean?